# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ 3 Preliminaries
Predictive Counterfactual Regret Minimization+ (PCFR+). PCFR+ (Farina et al., 2021) is a powerful variant of CFR+. It significantly outperforms CFR+ in many IIGs. PCFR+ employs Predictive RM+ (PRM+(Farina et al., 2021) as its local regret minimizer, with its key insight is to use the prediction. Specifically, as shown Figure 1, at each iteration $t$, PCFR+ maintains implicit and explicit accumulated counterfactual regrets: ${\bm{R}}^{t}_{I}$ and $\hat{\bm{R}}^{t-1}_{I}$. Firstly, PCFR+ makes a prediction and uses this prediction to derive new explicit accumulated counterfactual regrets $\hat{\bm{R}}^{t}_{I}$ from ${\bm{R}}^{t}_{I}$. Then, PCFR+ observes the instantaneous counterfactual regret $\bm{r}^{t}_{I}$ by following the strategy $\sigma^{t}$ defined by $\hat{\bm{R}}^{t}_{I}$. Lastly, $\bm{r}^{t}_{I}$ is subsequently used to derive ${\bm{R}}^{t+1}_{I}$ from ${\bm{R}}^{t}_{I}$. If the prediction aligns with the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$, Farina et al. (2021) show that the theoretical convergence of PCFR+ can be improved from $O(1/\sqrt{T})$ of CFR+ to $O(1/T)$. As tested in Farina et al. (2021), using the instantaneous counterfactual regret $\bm{r}^{t-1}_{I}$ observed at the previous iteration $t-1$ as the prediction is both simple and effective. Therefore, in practice, PCFR+ uses $\bm{r}^{t-1}_{I}$ as the prediction at iteration $t$. Formally, at each iteration $t$ and for each infoset $I\in\mathcal{I}$, PCFR+ updates its strategy according to
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t-1}_{I}]^{+},{\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (9)
@@
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and the second equality of the second line comes from the fact that $\hat{\bm{R}}^{t+1}_{I}\geq\bm{0}$.
4 Our Algorithms
PCFR+ leverages the prediction to accelerate the empirical convergence rate. However, when the prediction is inaccurate, its empirical convergence rate may decrease significantly, leading to unstable performance on certain IIGs. To enhance the robustness of PCFR+, in Section 4.1, we propose Asynchronous PCFR+ (APCFR+), which mitigates the impact of the prediction inaccuracy on convergence. We then provide a theoretical analysis for APCFR+ to demonstrate the reason why it enhances the robustness. Finally, we introduce Simple APCFR+ (SAPCFR+), a simple version of APCFR+, which requires only a single-line modification to the PCFR+ implementation, as detailed in Section 4.2. We show that SAPCFR+ achieves a constant-factor improvement in the theoretical convergence rate over PCFR+ in the worst case.
4.1 Asynchronou PCFR+ (APCFR+)
To mitigate the impact of the prediction inaccuracy on convergence of PCFR+, APCFR+ adaptively reduces the step-size when updating via the prediction, i.e., when updating the explicitly accumulated counterfactual regret. In other words, APCFR+ exploits the asynchrony of step-sizes between the updates of the implicit and explicit ones. Formally, at iteration $t$, the update rule of APCFR+ at infoset $I$ is
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}]^{+},{\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (10)
@@ 4.1 Asynchronou PCFR+ (APCFR+)
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and $\bm{r}^{0}_{I}=\bm{0}$. The comparison between the update rules of PCFR+ and APCFR+ has been shown in Figure 1. In the rest of this subsection, we first present the regret upper bound for APCFR+ with respect to any $\alpha^{t}_{I}$, as stated in Theorem 4.1. According to the discussion about Theorem 4.1, we show why APCFR+ can enhance the robustness of PCFR+ by mitigating the impact of the prediction inaccuracy on convergence. Lastly, we discuss how to automatically learn $\alpha^{t}_{I}$ from the regret bound shown in Theorem 4.1.
@@
$\footnotesize R^{T}(I)\leq\sqrt{{\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)}}.$
According to Theorem 4.1, APCFR+ mitigates the impact of the prediction inaccuracy, $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, on the regret upper bound (the theoretical convergence rate) by introducing the term $\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. This reduces the impact of the prediction inaccuracy on the regret upper bound compared to PCFR+ by a factor of 1/($1/(1+\alpha^{t}_{I})$ (note that when $\alpha^{t}_{I}=0$, APCFR+ reduces to PCFR+).
To assess the effectiveness of the asynchronization mechanism for step-sizes in decreasing the regret upper bound, we first present an example to show $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}\geq\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. Assume that, the number of actions at infoset $I$ is $2$, ${\bm{R}}^{t}_{I}$ is $[1;0]$, as well as $\bm{r}^{t-1}_{I}$ and $\bm{r}^{t}_{I}$ are $[1;-1]$ and $[-1;1]$, respectively. From the update rule of PCFR+, we can derive that ${\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+}=[0;1]$. Then, we have $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}=8>\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}=2$. This phenomenon can be explained by fact that the upper bound of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is only a quarter of the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, which is also used to derive SAPCFR+ (detailed in Section 4.2). In experiments, we also analyze the values of two terms $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}\ \text{and}\sum^{T}_{t=1}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2},$ for both PCFR+ and our algorithms (Figures 8 and 9). Among all algorithms, we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times than that of $\sum^{T}_{t=1}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. This indicates that introducing the term $\sum^{T}_{t=1}\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ and modifying the term $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ to $\sum^{T}_{t=1}\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}$, can reduce the regret upper bound. Furthermore, compared to PCFR+, both these two terms are smaller in our algorithms, further decreasing the regret upper bound. Then, we evaluate the values of ${\sum^{T}_{t=1}(\frac{\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2})},$ for both PCFR+ and our algorithms (Figures 10 and 11). In all games, the value of $\sum^{T}_{t=1}(\frac{\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2})$ is consistently smaller in our algorithms than in PCFR+. See more details and discussions in Appendix D.
Notably, Theorem 4.1 does not conflict the upper regret bound of CFR+ (where $\alpha^{t}_{I}\to\infty$), as it provides a larger upper regret bound than the original CFR+ upper bound. By altering the proof method, we get $R^{T}(I)\leq\sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}$ (detailed in Appendix B). By setting $\alpha^{t}_{I}\to\infty$, the original bound of CFR+ can be recovered. Additionally, for PCFR+ (where $\alpha^{t}_{I}\to 0$), the bound in Theorem 4.1 is identical to the one in its original version (the result in Theorem 3 of the original PCFR+ version can be easily improved to the bound presented in Theorem 4.1).
To eliminate the fine-tuning of $\alpha^{t}_{I}$, we propose an automated learning approach for $\alpha^{t}_{I}$. From Theorem 4.1, we have
To minimize the second line in the above equation, we can set $\alpha^{t}_{I}=\sqrt{{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}/{\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}}}.$ However, this is not feasible, as we need $\alpha^{t}_{I}$ to compute $\bm{r}^{t}_{I}$. Therefore, we adopt an alternative approach: at each iteration $t$,
[…]
Note that $5$ in Eq. (12) is included solely to ensure that the bound in Theorem 4.1 remains finite. In practice, we rarely observed $\alpha^{t}_{I}$ reaching 5 (Figures 6 and 7).
4.2 Simple APCFR+ (SAPCFR+)
We now introduce SAPCFR+, which is implemented with a single-line modification to the PCFR+ code. Specifically, SAPCFR+ sets $\alpha^{t}_{I}$ as a constant. To determine an appropriate $\alpha^{t}_{I}$, we use a value that ensures SAPCFR+ achieves a constant-factor reduction in the regret bound compared to PCFR+ in the worst case. Formally, as previously discussed, the key insight of finding such $\alpha^{t}_{I}$ lies in the fact that the upper bound of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is only a quarter of the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ (details are in the following). Formally, we first present Lemma 4.2.
@@ 4.2 Simple APCFR+ (SAPCFR+)
[Adapted from Lemma 11 of Wei et al. (2021)]. Assume that $T$ iterations of APCFR+ with any $\alpha^{t}_{I}\geq 0$ are conducted. Then for any infoset $I\in\mathcal{I}$ and $t\geq 1$, we have
$\footnotesize\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\leq\|{\bm{r}}^{t}_{I}\|^{2}_{2}.$
@@
Combining Theorem 4.1, Eq. (13), and Eq. (14), in the worst case, we obtain
@@
It is evident that when $\alpha^{t}_{I}=0$, i.e., for PCFR+, the worst-case counterfactual regret upper bound is given by
@@
Figure 3: Empirical convergence for CFR, CFR+, DCFR, PCFR+, Stable PCFR+, Smooth PCFR+, APCFR+, and SAPCFR+ in standard commonly used IIG benchmarks. Each algorithm runs for 5,000 iterations. In all plots,the x-axis is the number of iteration, and the y-axis represents exploitability,displayed on a logarithmic scale. Liar’s Dice ($x$) represents that every player is given a die with $x$ sides. Goofspiel ($x$) denotes that each player is dealt $x$ cards. Battleship ($x$) implies the size of grids is $x$.
When we set $\alpha^{t}_{I}=2$ (which comes from facts that (i) $2$ minimizes ${{\sum^{T}_{t=1}({4E}/{\alpha^{t}_{I}}+\alpha^{t}_{I}E)}}$ for any positive $E$ and (ii) ${{\sum^{T}_{t=1}({4E}/{(1+\alpha^{t}_{I})}+\alpha^{t}_{I}E)}}\leq{{\sum^{T}_{t=1}({4E}/{\alpha^{t}_{I}}+\alpha^{t}_{I}E)}}$), the counterfactual regret is bound by
@@
Clearly, setting $\alpha^{t}_{I}=2$ results in a lower regret upper bound. Therefore, for SAPCFR+, we set $\alpha^{t}_{I}=2$ for all $t\geq 1$. Formally, at each iteration $t$, SAPCFR+ updates its strategy at each infoset $I\in\mathcal{I}$ according to the following update rule:
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\frac{1}{3}\bm{r}^{t-1}_{I}]^{+},\ {\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (18)

### v2

@@ 3 Preliminaries
Predictive Counterfactual Regret Minimization+ (PCFR+). PCFR+ (farina2021faster) is a powerful CFR algorithm, which significantly outperforms other CFR algorithm in many IIGs. PCFR+ employs Predictive RM+ (PRM+(farina2021faster) as its local regret minimizer, with its key insight is to use the prediction. Specifically, as shown Figure 1, at each iteration $t$, PCFR+ maintains implicit and explicit accumulated counterfactual regrets: ${\bm{R}}^{t}_{I}$ and $\hat{\bm{R}}^{t-1}_{I}$. Firstly, PCFR+ makes a prediction and uses this prediction to derive new explicit accumulated counterfactual regrets $\hat{\bm{R}}^{t}_{I}$ from ${\bm{R}}^{t}_{I}$. Then, PCFR+ observes the instantaneous counterfactual regret $\bm{r}^{t}_{I}$ by following the strategy $\sigma^{t}$ defined by $\hat{\bm{R}}^{t}_{I}$. Lastly, $\bm{r}^{t}_{I}$ is subsequently used to derive ${\bm{R}}^{t+1}_{I}$ from ${\bm{R}}^{t}_{I}$. If the prediction aligns with the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$, farina2021faster show that the theoretical convergence of PCFR+ can be improved from $O(1/\sqrt{T})$ of CFR+ to $O(1/T)$. As tested in farina2021faster, using the instantaneous counterfactual regret $\bm{r}^{t-1}_{I}$ observed at the previous iteration t1t-1 as the prediction is both simple and effective. Therefore, in practice, PCFR+ uses $\bm{r}^{t-1}_{I}$ as the prediction at iteration $t$. Formally, at each iteration $t$ and for each infoset $I\in\mathcal{I}$, PCFR+ updates its strategy according to
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t-1}_{I}]^{+},\quad{\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (3)
@@
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and the forth equality comes from the fact that $\hat{\bm{R}}^{t}_{I}\geq\bm{0}$.
4 Methodology
PCFR+ leverages the prediction to accelerate the empirical convergence rate. However, when the prediction is inaccurate, its empirical convergence rate may decrease significantly, leading to unstable performance on certain IIGs. To enhance the robustness of PCFR+, we propose Asymmetric PCFR+ (APCFR+), which mitigates the impact of the prediction inaccuracy on the convergence rate via the adaptive asymmetry of step sizes. We then provide a theoretical analysis for APCFR+ to demonstrate the reason why it enhances the robustness. To simplify the implementation of APCFR+ due to the adaptive asymmetry, we propose Simple APCFR+ (SAPCFR+), using a constant asymmetry to guarantee that it can be implemented with a single-line modification compared to PCFR+.
4.1 Asymmetric PCFR+ (APCFR+)
To mitigate the impact of the prediction inaccuracy on convergence of PCFR+, APCFR+ adaptively reduces the step size when updating via the prediction, i.e., when updating the explicitly accumulated counterfactual regret. In other words, APCFR+ exploits the adaptive asymmetry of step sizes between the updates of the implicit and explicit ones. Formally, at iteration $t$ and infoset $I$, the update rule of APCFR+ is
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}]^{+},\quad{\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (4)
@@ 4.1 Asymmetric PCFR+ (APCFR+)
where $i=P(I)$, ${\bm{R}}^{1}_{I}=\bm{0}$, and $\bm{r}^{0}_{I}=\bm{0}$. The comparison between the update rules of PCFR+ and APCFR+ has been shown in Figure 1. In the rest of this subsection, we first present the regret upper bound for APCFR+ with respect to any $\alpha^{t}_{I}$, as stated in Theorem 4.1. According to the discussion about Theorem 4.1, we show why APCFR+ can enhance the robustness of PCFR+ by mitigating the impact of the prediction inaccuracy on the convergence rate. Lastly, we discuss how to automatically learn $\alpha^{t}_{I}$ from the regret bound shown in Theorem 4.1.
@@
$R^{T}(I)\leq\sqrt{{\sum^{T}_{t=1}\left(\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\right)}}.$ (5)
Why the asymmetry mechanism is effective. To assess the effectiveness of the asymmetry mechanism for step sizes in decreasing the regret upper bound (improving the convergence rate), we show the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is four times than that of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. Firstly, we introduce Lemma 4.2.
@@
[Adapted from Lemma 11 of wei2020linear]. Assume that $T$ iterations of APCFR+ with any $\alpha^{t}_{I}\geq 0$ are conducted. Then for any infoset $I\in\mathcal{I}$ and $t\geq 1$, we have
$\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}\leq\|{\bm{r}}^{t}_{I}\|^{2}_{2}.$ (6)
@@
In experiments, we also analyze the values of two terms $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}\ \text{and}\sum^{T}_{t=1}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2},$ for both PCFR+ and our algorithms (Figures 6 and 7). Among all algorithms, we observe that the value of $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ is at least three times than that of $\sum^{T}_{t=1}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$. This indicates that introducing the term $\sum^{T}_{t=1}\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ and modifying the term $\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$ to $\sum^{T}_{t=1}\frac{\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}$, can reduce the regret upper bound. Furthermore, compared to PCFR+, both of these two terms are smaller in our algorithms, further decreasing the regret upper bound. Then, we evaluate the values of $\sum^{T}_{t=1}(\frac{\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2})$}, for both PCFR+ and our algorithms (Figures 8 and 9). In all games, the value of $\sum^{T}_{t=1}(\frac{\|{\bm{r}}^{t}_{I}-{\bm{r}}^{t-1}_{I}\|^{2}_{2}}{1+\alpha^{t}_{I}}+\alpha^{t}_{I}\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2})$ is consistently smaller in our algorithms than in PCFR+. See more details in Appendix D.
An alternative regret upper bound of APCFR+. Notably, Theorem 4.1 does not conflict the upper regret bound of CFR+ (where $\alpha^{t}_{I}\to\infty$), as it provides a larger upper regret bound than the original CFR+ upper bound. By altering the proof method, we get $R^{T}(I)\leq\sqrt{\sum^{T}_{t=1}\|{\bm{r}}^{t}_{I}-\frac{1}{1+\alpha^{t}_{I}}\bm{r}^{t-1}_{I}\|^{2}_{2}}$, as shown in Theorem B.1 (detailed in Appendix B). By setting $\alpha^{t}_{I}\to\infty$, the original bound of CFR+ can be recovered. Additionally, for PCFR+ (where $\alpha^{t}_{I}\to 0$), the bound in Theorem 4.1 is identical to the one in its original version (the result in Theorem 3 of the original PCFR+ version can be easily improved to the bound presented in Theorem 4.1). The reason why we employ Theorem 4.1 in the main text rather than Theorem B.1 is that the regret bound in Theorem B.1 is typically larger than that in Theorem 4.1, as demonstrated in Appendix D (Figures 8, 9, 10, and 11).
Automatic learning approach for $\alpha^{t}_{I}$. To eliminate the fine-tuning of $\alpha^{t}_{I}$, we propose an automatic learning approach for $\alpha^{t}_{I}$. From Theorem 4.1, we have
To minimize the right-hand side of the last inequality, we can set $\alpha^{t}_{I}=\sqrt{\frac{\|\bm{r}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}}{\|\bm{R}^{t+1}_{I}-\bm{R}^{t}_{I}\|^{2}_{2}}}.$ However, this is not feasible, as we need $\alpha^{t}_{I}$ to compute $\bm{r}^{t}_{I}$. Therefore, we adopt an alternative approach:
$\alpha^{t}_{I}=\min\left(\sqrt{\frac{\sum_{\tau=1}^{t-1}\|\bm{r}^{\tau}_{I}-\bm{r}^{\tau-1}_{I}\|^{2}_{2}}{\sum_{\tau=1}^{t-1}\|\bm{R}^{\tau+1}_{I}-\bm{R}^{\tau}_{I}\|^{2}_{2}}},\alpha_{max}\right).$ (10)
Note that the parameter in $\alpha_{max}$ in Eq. (10) is included solely to ensure that the bound in Theorem 4.1 remains finite. In this paper, we directly set it as $5$ to reduce the cost of hyperparameter tuning. In practice, we rarely observed $\alpha^{t}_{I}$ reaching 5 (Figures 4 and 5).
4.2 Simple APCFR+ (SAPCFR+)
To simplify the implementation of APCFR+ caused by the automatic learning approach of $\alpha^{t}$, we introduce SAPCFR+, which is implemented with a single-line modification to the PCFR+ code. Specifically, SAPCFR+ sets $\alpha^{t}_{I}=2$. The key insight of setting $\alpha^{t}_{I}=2$ lies in the fact that the upper bound of $\|{\bm{R}}^{t+1}_{I}-{\bm{R}}^{t}_{I}\|^{2}_{2}$ is only a quarter of the upper bound of $\|{\bm{r}}^{t}_{I}-\bm{r}^{t-1}_{I}\|^{2}_{2}$, as shown in Eq. (7), and Eq. (8).
Specifically, combining Theorem 4.1, Eq. (7), and Eq. (8), in the worst case, we obtain
@@ 4.2 Simple APCFR+ (SAPCFR+)
It is evident that when $\alpha^{t}_{I}=0$, i.e., for PCFR+, the worst-case counterfactual regret upper bound is
@@
From the facts that (i) $2$ minimizes $\sum^{T}_{t=1}({4E}/{\alpha^{t}_{I}}+\alpha^{t}_{I}E)$}} for any positive $E$ and (ii) ${{\sum^{T}_{t=1}({4E}/{(1+\alpha^{t}_{I})}+\alpha^{t}_{I}E)}}\leq{{\sum^{T}_{t=1}({4E}/{\alpha^{t}_{I}}+\alpha^{t}_{I}E)}}$, we can set $\alpha^{t}_{I}=2$, which implies the counterfactual regret is bound by
@@
Clearly, setting $\alpha^{t}_{I}=2$ results in a lower regret upper bound than PCFR+. Therefore, for SAPCFR+, we set $\alpha^{t}_{I}=2$ for all $t\geq 1$. Formally, at each iteration $t$, SAPCFR+ updates its strategy at each infoset $I\in\mathcal{I}$ according to the following update rule:
$\hat{\bm{R}}^{t}_{I}=[{\bm{R}}^{t}_{I}+\frac{1}{3}\bm{r}^{t-1}_{I}]^{+},\quad{\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+},$ (14)